OpenStudy (anonymous):

Find the sum of the first 30 positive multiples of 3.

OpenStudy (anonymous):

arithmetic series

OpenStudy (anonymous):

a = 3 d=3 n=30

OpenStudy (curry):

i believe its 1395

OpenStudy (curry):

not sure check just in case

OpenStudy (curry):

jimmy is it right?

OpenStudy (anonymous):

use Sn = n/2 [2a + (n-1)d]

OpenStudy (anonymous):

ill check

OpenStudy (curry):

i thought the equation is Sn = n/2(A1+An)

OpenStudy (curry):

i guess its a variation

OpenStudy (anonymous):

15 *( 6 + 29*3) = 15 * 93 = 1395

OpenStudy (curry):

yep its was 1395

OpenStudy (curry):

so jimmy Sn = n/2(A1+An) also works

OpenStudy (curry):

its the one i was taught

OpenStudy (curry):

faster to do

OpenStudy (anonymous):

yes - when An = last term

OpenStudy (curry):

yep